OpenStudy (anonymous):

Express in interval notation: I5x+2I <13

OpenStudy (anonymous):

\[|5x+2<13|\] \[-13<5x+2<13\] \[-15<5x<11\] \[-3<x<\frac{11}{5}\]

OpenStudy (anonymous):

interval notation \[(-3,\frac{11}{5})\]

OpenStudy (anonymous):

depending on where you come from. could be \[]-3,\frac{11}{5}[\]

OpenStudy (anonymous):

Thank you!